### v1

# Convective Flows in Sheared Packings of Spherical Particles

Abstract
Understanding how granular materials respond to shear stress remains a central challenge in soft matter physics. We report direct observations of persistent granular convection in the bulk shear zones of spherical particle packings— a phenomenon previously associated primarily with particle shape anisotropy or boundary effects. By employing various bead-coloring techniques in a split-bottom geometry, we reveal internal flow fields within sheared granular packings. We find robust convective structures, strikingly governed by system geometry: at low filling heights, two counter-rotating convection rolls emerge, while at higher filling heights, a single dominant convection cell forms, featuring radially outward flow at the surface. This transition is driven by the height-dependent broadening of the shear zone, which introduces shear rate asymmetry across its flanks. Notably, the transition occurs entirely within the open shear band regime. These findings demonstrate the crucial role of system geometry for secondary flow formation in dense packings of frictional particles, with significant implications for geophysical flows and industrial processes.
preprint: AIP/123-QED
Granular flow under shear presents fundamental challenges with wide-ranging implications, from powder processing to geological faulting. In slowly sheared granular systems, shear strain localizes into shear zones— often narrow regions near boundaries, but also broad zones within the bulk Schall and van Hecke (2010); Mueth et al. (2000); Kollmer et al. (2020); Fenistein and van Hecke (2003); Shaebani et al. (2021); Cheng et al. (2006); Unger et al. (2004); Török et al. (2007); Moosavi et al. (2013); Henann and Kamrin (2013); Artoni et al. (2018). While such primary, shear-aligned flows have been extensively characterized, the emergence of secondary flows— circulatory motions transverse to the main shear direction— remains less understood. Although typically weaker than the primary flow, secondary convective flows can strongly influence bulk transport, particle mixing, and segregation Harrington et al. (2013); Fan and Hill (2010); Khosropour et al. (1997), making them crucial in both natural and industrial contexts.
Convective rolls in granular flows are often attributed to narrow shear zones near boundaries, such as in Couette geometries, where the interplay of shear-induced dilatancy, gravity, and centrifugal forces drives circulatory motion in radial and vertical directions Murdoch et al. (2013); Krishnaraj and Nott (2016); Zheng et al. (2020); Khosropour et al. (1997). However, many natural and industrial flows involve spatially extended shear zones that are not confined to boundaries. This raises a key question: Is the generation of secondary flows a generic feature of sheared granular matter, or does it fundamentally require boundaries? Split-bottom geometries— designed to probe wide shear zones away from lateral walls Fenistein and van Hecke (2003); Shaebani et al. (2021); Cheng et al. (2006); Unger et al. (2004); Török et al. (2007); Moosavi et al. (2013); Henann and Kamrin (2013); Harrington et al. (2013)— provide a powerful platform to address this. In such systems, convective flows have been observed in binary mixtures Harrington et al. (2013); Fan and Hill (2010), and surface heaping accompanied by secondary flows was reported for highly elongated particles Fischer et al. (2016); Wortel et al. (2015), where misalignment between particle orientation and flow streamlines was proposed as the driving mechanism. In contrast, spherical particles exhibit no surface heaping, reinforcing the prevailing assumption that significant secondary convection is absent. While recent simulations suggest possible secondary motion in split-bottom geometries Dsouza and Nott (2021), direct experimental evidence for persistent secondary flows in spherical particle packings far from boundaries has remained elusive.
Figure 1: Schematic (left) and photograph (right) of the split-bottom setup used in the experiments.
Figure 2: (a) Schematic of initial vertical plane of colored tracers embedded along one diameter of the cylinder and initial surface photo before shearing. (b) Deformed tracer line after shear reveals local angular displacement at different bulk heights $h$. (c) Angular velocity profiles $\omega(r)$ at various $h$ (solid lines) and corresponding error function fits (dashed lines). Inset: axial angular velocity $\omega_{0}$ vs $h$. The line represents a Gaussian fit. (d) Comparison of shear zone center $R_{c}$ and width $\delta$ in our setup (green) with prior rough-boundary data (purple). Solid lines and symbols represent $R_{c}$ and shaded regions indicate $\delta$; see text. (e) Variational model prediction of $R_{c}$ versus $\mu_{\text{rel}}$ for different $H$. (f) $\delta$ vs $h$ for different values of $H$. The dotted line represents $\delta{\sim}h^{0.58}$.
Here, we experimentally reveal that dense packings of spherical particles under shear can exhibit sustained bulk convection. Using a split-bottom cell combined with bead-coloring techniques, we visualize internal flow fields and uncover robust convective structures deep within the bulk— despite the absence of significant surface heaping, consistent with previous surface-based observations Fischer et al. (2016); Sakaie et al. (2008); Cabrera and Polania (2020); Mohammadi et al. (2022). Strikingly, we observe a geometry-controlled transition in convection patterns from counter-rotating rolls at low filling heights to a single dominant convection cell at higher heights. We show that this transition is driven by the filling-height-dependent broadening of the shear zone, which induces asymmetry in local shear rates across the flow. These findings demonstrate that even in the absence of particle anisotropy or boundaries, geometry alone can generate persistent secondary flows in spherical granular systems, offering fresh insights into shear-induced transport in dense granular media.
Setup and procedure.— Our experiments use a split-bottom geometry consisting of a rotating disk of radius $R_{s}{=}7.5\,\text{cm}$ flush with a fixed outer ring inside a cylindrical cell of radius $R{=}10\,\text{cm}$ (Fig. 1). The disk is rotated at a constant angular velocity $\Omega{=}0.127\,\text{rad/s}$. The granular medium comprises $2\,\text{mm}$ glass beads with 15% polydispersity, which has negligible effect on packing fabric Shaebani et al. (2012), and is filled to height $H$ with a bulk packing fraction of  ${\sim}\,60\%$. Colored tracer grains are used to visualize internal flow. A standardized pouring protocol, with gently leveling the surface every  ${\sim}\,8\,\text{cm}$, ensures reproducibility. Smooth chamber walls allow axial slip, promoting shear localization within the bulk. To reconstruct internal flow patterns, we photograph horizontal cross-sections from above after carefully removing successive top layers, yielding 2D tomographic maps of tracer positions within the bulk.
Bulk primary flow profiles.— To characterize the primary flow field, we first determine the bulk shear profile. To this aim, we design the initial conditions such that a vertical plane of black-colored tracer particles is embedded along one diameter of the cylinder, forming a rectangular sheet from the bottom to the free surface and spanning the full cylinder width [Fig. 2(a)]. After shearing the system for $210\,\text{s}$, we photograph the surface, then sequentially remove  ${\sim}\,2\,\text{mm}$ slices from the top and image each exposed layer. This procedure yields the deformed shape of the initially straight tracer line at each bulk height $h$ [Fig. 2(b)], from which local angular displacements and velocity profiles $\omega(r)$ at various depths can be reconstructed; see Fig. 2(c) and Refs. Fenistein and van Hecke (2003); Shaebani et al. (2021). These profiles are well described by an error function $\omega{=}\frac{\omega_{0}}{2}\left[1{-}\text{erf}\left(\frac{r{-}R_{c}}{\delta}\right)\right]$, allowing extraction of the axial angular velocity $\omega_{0}$ (inset of panel c) and the center position $R_{c}$ and width $\delta$ of the shear zone (panel d). Notably, $\omega_{0}$ is small even near the rotating base due to slip at the bottom boundary, and decays rapidly with $h$, following a Gaussian trend Cheng et al. (2006); Madani et al. (2021).
Figure 3: (a) Top-view images at selected time points for various filling heights $H$, showing the dispersion of colored tracer grains initially placed in narrow radial bands on the surface. (b) Mean radial position $\langle r\rangle$ of tracers as a function of time, for different initial radii and filling heights. Changes in $\langle r\rangle$ reflect surface signatures of underlying convective flow.
Figure 4: (a) Initial configuration of colored tracers placed at mid-height (h=H/2h{=}H{/}2) for two filling levels, covered with a black tracer layer at the top. (b) Reconstructed bulk tracer positions after shear for $H{=}4\,\text{cm}$. (c) Tracer density map in the ($r,h$) plane for $H{=}4\,\text{cm}$ in the steady state, overlaid with the shear zone center and boundaries, indicating a single dominant convective roll. Inset: Side view of the initial arrangement of colored tracers.
Figure 5: (a) Reconstructed bulk tracer positions after shear for $H{=}2\,\text{cm}$. (b) Individual tracer density maps for $H{=}2\,\text{cm}$ in the steady state, overlaid with the shear zone center and boundaries, indicating two counter-rotating convective rolls. Inset: Side view of the initial arrangement of colored tracers. (c) Tracer density maps for $H{=}2\,\text{cm}$ and bulk heights $h{=}0.5,\,1.0,\,2.0\,\text{cm}$ at an early time point $t\,{=}\,3\,\text{min}$.
Comparing $R_{c}$ and $\delta$ in our setup with known results in similar but rough-bottom geometries— where the behavior is well captured by $h{=}H{-}R_{c}\big[1{-}\frac{R_{s}}{R_{c}}\big(1{-}(\frac{H}{R_{s}})^{5/2}\big)\big]^{2/5}$ and $\delta{\sim}h^{\beta}$ ($0.5{<}\beta{<}1$) Fenistein and van Hecke (2003); Unger et al. (2004); Török et al. (2007)— reveals significant deviations in $R_{c}$ and $\delta$ due to slip [Fig. 2(d)]. We confirm this through numerical minimization of energy dissipation using a variational approach Unger et al. (2004); Török et al. (2007); Shaebani et al. (2021); Madani et al. (2021); Unger (2007), showing that reduced bottom friction— characterized by an effective friction $\mu_{\text{rel}}{=}\mu_{\text{bottom}}^{\text{eff}}{/}\mu_{\text{bulk}}^{\text{eff}}$— shifts $R_{c}$ inward [Fig. 2(e)]. As shown in Fig. 2(f), shear zone broadening due to slip enhances with filling height— which increases shear rate asymmetry across the flanks, an essential ingredient for the emergence of convection. These results justify our choice of a smooth-bottom setup to access broader shear zones and thereby amplify the secondary flow signatures.
Surface radial flow.— To probe secondary flow near the surface, we place colored tracer grains (red, blue, green, black) in narrow radial bands at different radii atop the transparent bed [Fig. 3(a)] and track their radial displacement over time. The mean tracer radius $\langle r\rangle$ for each color is monitored at 30-minute intervals for various filling heights $H{=}1{-}4\,\text{cm}$. At $H{=}1\,\text{cm}$, only black tracers near the outer edge ($r{=}85\,\text{mm}$) exhibit slight spreading and other colors remain stationary, indicating negligible surface convection [Fig. 3(b)].
At $H{=}2\,\text{cm}$, tracers at radii $r{=}50,\,75\,\text{mm}$ (near the flanks of the surface shear zone) disperse and converge toward an intermediate radius, while inner tracers at $r{=}15,\,30\,\text{mm}$ remain fixed. This convergence indicates the presence of two counter-rotating convective rolls, featuring either downward flow along the shear flanks and upward return flow between them, or the reverse pattern. During the initial stage (up to $t=3~\mathrm{min}$), black particles at $r=75~\mathrm{mm}$ move outward, while green particles at $r=50~\mathrm{mm}$ move inward. Subsequently, the particles in this region descend through the convective rolls and re-emerge at the surface at radii different from their initial positions, leading to a shift in their mean radii in the opposite direction [Fig. 3(b)]. The initial divergence and subsequent convergence of the black and green particles indicate a downward flow along the shear flanks and an upward return flow between them.
For $H{=}3\,\text{cm}$, the surface flow is strongest: tracers at $r{=}30,\,50\,\text{mm}$ move outward up to radial distance $r{\approx}70\,\text{mm}$, consistent with a dominant single convection roll with downward flow at the outer shear flank at $r{\simeq}70\,\text{mm}$ and upward return flow at the inner shear flank at $r{\simeq}30\,\text{mm}$. Tracers at $r{=}15,\,75\,\text{mm}$ show minimal displacement, indicating they lie outside the active convective region.
At $H{=}4\,\text{cm}$, surface flow is largely suppressed, though there still exists a weak radial motion. The peak radial surface velocity $v_{r}\,{\approx}\,6{\times}10^{-4}\,\text{mm/s}$ is about 2% of the peak azimuthal speed $v_{\theta}\,{\approx}\,3{\times}10^{-2}\,\text{mm/s}$, reflecting the dynamical relevance of convective flows at this height.
Figure 6: (a) Initial configuration of horizontal red, green, and blue tracer layers at depths $h{=}0.5,\,2,\,3\,\text{cm}$, respectively, in a granular pile of height $H{=}4\,\text{cm}$. (b) Reconstructed positions of the tracers after 160 minutes of shear. (c) 2D tracer density map in the ($r,h$) plane, showing radial and vertical displacements. Shear band center and boundaries are overlaid. Inset: Side view of the initial arrangement of colored tracers.
Bulk convective flow.— To probe internal convection, we prepare a single colored layer at mid-depth (h=H/2h{=}H{/}2) for two filling heights $H{=}2,\,4\,\text{cm}$. Tracers are arranged in three radial bands: red beads in the inner region ($r{<}6\,\text{cm}$), green in a ring ($6{<}r{<}7\,\text{cm}$), and blue in the outer bulk ($7{<}r{<}10\,\text{cm}$), with a top layer of black tracers covering the free surface [Fig. 4(a)]. After shearing the system for 60 min ($H{=}2\,\text{cm}$) or 180 min ($H{=}4\,\text{cm}$), the system is excavated layer by layer, and the bulk tracer positions are reconstructed [Figs. 4(b) and 5(a)]. For $H{=}4\,\text{cm}$, the tracer density map in the ($r,h$) plane shown in Fig. 4(c) reveals a strong single convection roll: red tracers rise while green and blue tracers descend within the shear zone. Black surface tracers are displaced outward and drawn into the bulk at the outer flank. A small surface elevation in the central region suggests dilatancy induced by flow. The entire convective motion is confined within the shear zone, with negligible transport beyond its boundaries.
For $H{=}2\,\text{cm}$, tracer trajectories [Fig. 5(b)] reveal that the flow direction of the two counter-rotating convection rolls is downward at the shear flanks: Tracers from all colors descend along bulk shear flanks, reappear at the top free surface, and move toward the edges of the shear zone, consistent with the surface flow observations. Black surface tracers are drawn toward the surface shear zone edges and disappear into the bulk, further corroborating the dual-roll structure. Formation of two counter-rotating convection rolls for $H{=}2\,\text{cm}$ is also evident when we track the colored beads at an earlier time point $t\,{=}\,3\,\text{min}$ for bulk heights $h{=}0.5,\,1.0,\,2.0\,\text{cm}$; see Fig. 5(c). Overall, convective transport is localized within the shear zone and highly sensitive to the filling height.
To further improve the bulk resolution at large filling height $H{=}4\,\text{cm}$, we implement a layered tracer configuration. Colored grains are arranged in horizontal planes: red at $h{=}0.5\,\text{cm}$, green at $h{=}2\,\text{cm}$, and blue at $h{=}3\,\text{cm}$ [Fig. 6(a)]. After shearing the system for 160 mins, we excavate the packing layer by layer, capturing the tracer positions at each step [Fig. 6(b)]. This procedure enables us to reconstruct a 2D density map of tracer positions in the ($r,h$) plane [Fig. 6(c)]. The tracer displacements reveal clear radial and vertical motions indicative of secondary flow in the bulk. Red particles near the base migrate inward and then upward, tracing a clockwise circulation. Green tracers from the mid-plane predominantly descend and accumulate near the bottom within the shear zone, while a portion at smaller radii is lifted upward— again consistent with toroidal motion. Blue tracers, initially near the top, move downward at larger radii. The overall pattern confirms the presence of a dominant single convection roll confined mainly within the bulk shear zone.
Dilatancy and heap formation.— To examine possible surface deformation and dilatancy during shear, we employ a green laser sheet (532 nm, Fath Optics) to measure changes in the free surface profile before and after rotation. Because the transparent glass beads scatter light and broaden the laser sheet, we apply a thin opaque ash layer (thickness < $75\,\mu\text{m}$) to the surface to improve optical contrast. The laser is projected at an oblique angle, and images of the illuminated surface are captured from a fixed viewing angle. The surface profiles are extracted using custom MATLAB scripts. Figure 7(a) shows the laser sheet projected on the ash-covered surface. Profiles are recorded along six diameters (panel b), and the azimuthally averaged height change relative to the initial flat surface is plotted in panel c. A slight surface elevation is observed primarily in the region $1{<}r{<}4\,\text{cm}$. The estimated post-shear packing fraction is approximately 58%, reduced from the initial 60%, indicating modest dilatancy. The maximum heap height remains small— less than 3% of the filling height— much lower than the prominent heap formation (${\sim}20\%$) reported for systems of elongated particles Fischer et al. (2016); Wortel et al. (2015). Thus, the absence of pronounced heap formation at the surface does not preclude the presence of bulk convection.
Onset of convection roll transition.— To rationalize the transition from two counter-rotating convection rolls at low filling heights to a single dominant roll at higher fillings, we consider the radial shear rate distribution at the free surface implied by the angular velocity profile. Assuming the error function $\omega(r)$ for the angular velocity profile, the associated tangential velocity can be obtained as $v_{\theta}(r){=}\,r\,\omega(r)$, from the radial derivative of which, the shear rate can be obtained as $\dot{\gamma}(r){=}\frac{1}{2}(\partial_{r}v_{\theta}{-}v_{\theta}{/}r)=\frac{1}{2}r\partial_{r}\omega$. We estimate two maxima of the shear rate by computing the shear rates at the flanks of the shear zone (defined as $R_{c}{\pm}\delta$). Their ratio $\dot{\gamma}_{\text{out}}{/}\dot{\gamma}_{\text{in}}$ is plotted in Fig. 7(d) as a function of $H{\propto}\,\delta^{1{/}\beta}$ ($\beta{=}0.5$). This ratio increases monotonically with increasing $H$, reflecting growing asymmetry in shear strength. Note that the strength of shearing directly influences the local dilatancy and, thus, the strength of gravity-induced downward convective roll. At small $H$, the shear rates— and consequently the downward driving strength for convective motion— are comparable at both flanks, supporting coexistence of counter-rotating rolls. For wider shear zones, the outer flank experiences significantly higher shear, making its associated clockwise convective roll dominant. This mechanism explains the experimentally observed crossover to a single-roll regime at larger filling heights. It also suggests that a similar transition could occur deep within tall granular piles. However, gravity-induced confining stresses at large depths suppress particle rearrangements Shaebani et al. (2007, 2008) and thus convective roll formation— a constraint that can be lifted under reduced gravity conditions.
Figure 7: (a) Laser-sheet visualization of the free surface after shear, with ash applied for contrast enhancement. (b) Extracted surface profiles along six diameters across the pile. (c) Azimuthally averaged surface height change (relative to the initial flat surface), showing slight elevation toward the cell center. (d) Fraction of shear rates at the outer and inner flanks of the surface shear zone vs $H$. The asymmetry grows with filling height, promoting a transition from symmetric counter-rotating convection rolls to a dominant single roll as shear becomes localized more strongly at the outer edge.
We have demonstrated direct experimental evidence of geometry-driven convective flow in sheared granular packings of spherical particles. Using tracer layers and tomographic reconstruction, we reveal a transition from symmetric counter-rotating convection rolls at low filling heights to a dominant single-roll pattern at higher fillings. This transition correlates with increasing asymmetry in the radial shear profile, consistent with a shear-driven convection mechanism. These findings highlight the intrinsic ability of granular systems to sustain bulk transport through internal flow structures, even in the absence of particle anisotropy or boundary effects. Such geometry-induced convection can play a crucial role in mixing, segregation, and particle migration in granular media, with direct relevance to natural processes and industrial applications.
DATA AVAILABILITY
The data that support the findings of this study are available from the corresponding author upon reasonable request.