### v2

# Convective Flows in Sheared Packings of Spherical Particles

Abstract
Understanding how granular materials respond to shear stress remains a central challenge in soft matter physics. We report direct observations of persistent granular convection in the bulk shear zones of spherical particle packings— a phenomenon previously associated primarily with particle shape anisotropy or boundary effects. By employing various bead-coloring techniques in a split-bottom geometry, we reveal internal flow fields within sheared granular packings. We find robust convection rolls, strikingly governed by system geometry: at low filling heights, two counter-rotating convection rolls emerge, while at higher filling heights, a single dominant convection roll forms, featuring radially outward flow at the surface. This transition is driven by the height-dependent broadening of the shear zone, which introduces shear rate asymmetry across its flanks. Notably, the transition occurs entirely within the open shear band regime. These findings underscore the pivotal role of system geometry in shaping secondary flow formation in dense packings of frictional particles, suggesting possible broader relevance to geophysical flow dynamics and industrial applications.
preprint: AIP/123-QED
Granular flow under shear presents fundamental challenges with wide-ranging implications, from powder processing to geological faulting. In slowly sheared granular systems, shear strain localizes into shear zones— often narrow regions near boundaries, but also broad zones within the bulk Schall and van Hecke (2010); Mueth et al. (2000); Kollmer et al. (2020); Fenistein and van Hecke (2003); Shaebani et al. (2021); Cheng et al. (2006); Unger et al. (2004); Török et al. (2007); Moosavi et al. (2013); Henann and Kamrin (2013); Artoni et al. (2018). While such primary, shear-aligned flows have been extensively characterized, the emergence of secondary flows— circulatory motions transverse to the main shear direction— remains less understood. Although typically weaker than the primary flow, secondary convective flows can strongly influence bulk transport, particle mixing, and segregation Harrington, Weijs, and Losert (2013); Fan and Hill (2010); Khosropour et al. (1997), making them crucial in both natural and industrial contexts.
Convection rolls in granular flows are often attributed to narrow shear zones near boundaries, such as in Couette geometries, where the interplay of shear-induced dilatancy, gravity, and centrifugal forces drives circulatory motion in radial and vertical directions Murdoch et al. (2013); Krishnaraj and Nott (2016); Zheng, Luo, and Yu (2020); Khosropour et al. (1997). However, many natural and industrial flows involve spatially extended shear zones that are not confined to boundaries. This raises a key question: Is the generation of secondary flows a generic feature of sheared granular matter, or does it fundamentally require boundaries? Split-bottom geometries— designed to probe wide shear zones away from lateral walls Fenistein and van Hecke (2003); Shaebani et al. (2021); Cheng et al. (2006); Unger et al. (2004); Török et al. (2007); Moosavi et al. (2013); Henann and Kamrin (2013); Harrington, Weijs, and Losert (2013)— provide a powerful platform to address this. In such systems, convective flows have been observed in binary mixtures Harrington, Weijs, and Losert (2013); Fan and Hill (2010), and surface heaping accompanied by secondary flows was reported for highly elongated particles Fischer et al. (2016); Wortel et al. (2015), where misalignment between particle orientation and flow streamlines was proposed as the driving mechanism. In contrast, spherical particles exhibit no surface heaping, reinforcing the prevailing assumption that significant secondary convection is absent. While recent simulations suggest possible secondary motion in split-bottom geometries Dsouza and Nott (2021), direct experimental evidence for persistent secondary flows in spherical particle packings far from boundaries has remained elusive.
Figure 1: Schematic (left) and photograph (right) of the split-bottom setup used in the experiments.
Figure 2: (a) Schematic of initial vertical plane of colored tracers embedded along one diameter of the cylinder and initial surface photo before shearing. (b) Deformed tracer line after shear reveals local angular displacement at different bulk heights $h$. (c) Angular velocity profiles $\omega(r)$ at various $h$ (solid lines) and corresponding error function fits (dashed lines). Inset: axial angular velocity $\omega_{0}$ vs $h$. The line represents a Gaussian fit. (d) Comparison of shear zone center $R_{c}$ and width $\delta$ in our setup (green) with prior rough-boundary data (purple). Solid lines and symbols represent $R_{c}$ and shaded regions indicate $\delta$; see text. (e) Variational model prediction of $R_{c}$ versus $\mu_{\text{rel}}$ for different $H$. (f) $\delta$ vs $h$ for different values of $H$. The dotted line represents $\delta{\sim}h^{0.58}$.
Here, we experimentally reveal that dense packings of spherical particles under shear can exhibit sustained bulk convection. Using a split-bottom cell combined with bead-coloring techniques, we visualize internal flow fields and uncover robust convection rolls deep within the bulk— despite the absence of significant surface heaping, consistent with previous surface-based observations Fischer et al. (2016); Sakaie et al. (2008); Cabrera and Polania (2020); Mohammadi et al. (2022). Strikingly, we observe a geometry-controlled transition in convection patterns from counter-rotating rolls at low filling heights to a single dominant convection cell at higher heights. We show that this transition is driven by the filling-height-dependent broadening of the shear zone, which induces asymmetry in local shear rates across the flow. These findings demonstrate that even in the absence of particle anisotropy or boundaries, geometry alone can generate persistent secondary flows in spherical granular systems, offering fresh insights into shear-induced transport in dense granular media.
Figure 3: (a) Top-view images at selected time points for various filling heights $H$, showing the dispersion of colored tracer grains initially placed in narrow radial bands on the surface. (b) Mean radial position $\langle r\rangle$ of tracers as a function of time, for different initial radii and filling heights. Changes in $\langle r\rangle$ reflect surface signatures of underlying convective flow.
Setup and procedure.— Our experiments use a split-bottom geometry consisting of a rotating disk of radius $R_{s}{=}7.5\,\text{cm}$ flush with a fixed outer ring inside a cylindrical cell of radius $R{=}10\,\text{cm}$ (Fig. 1). The disk is rotated at a constant angular velocity $\Omega{=}0.127\,\text{rad/s}$ ( ${\simeq}\,1.213\,\text{rpm}$). The angular velocity of the disk is chosen sufficiently low to ensure operation in the slow quasistatic granular‐flow regime, in which the flow patterns are known to be essentially shear‐rate independent MiDi (2004); Fenistein, van de Meent, and van Hecke (2004). The granular medium comprises $2\,\text{mm}$ glass beads with 15% polydispersity, which has negligible effect on packing fabric Shaebani et al. (2012). Colored tracer grains are used to visualize internal flow. All experiments employing colored tracers use the same base glass beads. The applied color coating adds only a few micrometers to the bead diameter, which is negligible compared with their size. The cylinder is filled to height $H$ with a bulk packing fraction of  ${\sim}\,60\%$. A standardized pouring protocol, with gently leveling the surface every  ${\sim}\,8\,\text{cm}$, ensures reproducibility. Smooth chamber walls allow axial slip, promoting shear localization within the bulk. To reconstruct internal flow patterns, we photograph horizontal cross-sections from above after carefully removing successive top layers, yielding 2D tomographic maps of tracer positions within the bulk. We note that, despite the relatively large bead size, random packing readily produces smooth shear zones and desired filling heights with less than $1\%$ variation across the large cylinder. The presence of many beads in the shear zone and surface layer smooths surface fluctuations and minimizes finite-size effects. Using significantly smaller beads is not practical, as electrostatic charging and cohesive forces become relevant, causing the material to deviate from dry, cohesionless granular behavior.
Bulk primary flow profiles.— To characterize the primary flow field, we first determine the bulk shear profile. To this aim, we design the initial conditions such that a vertical plane of black-colored tracer particles is embedded along one diameter of the cylinder, forming a rectangular sheet from the bottom to the free surface and spanning the full cylinder width [Fig. 2(a)]. After shearing the system for $210\,\text{s}$, we photograph the surface, then sequentially remove  ${\sim}\,2\,\text{mm}$ slices from the top and image each exposed layer. This procedure yields the deformed shape of the initially straight tracer line at each bulk height $h$ [Fig. 2(b)], from which local angular displacements and velocity profiles $\omega(r)$ at various depths can be reconstructed; see Fig. 2(c) and Refs. Fenistein and van Hecke (2003); Shaebani et al. (2021). These profiles are well described by an error function $\omega{=}\frac{\omega_{0}}{2}\left[1{-}\text{erf}\left(\frac{r{-}R_{c}}{\delta}\right)\right]$, allowing extraction of the axial angular velocity $\omega_{0}$ (inset of panel c) and the center position $R_{c}$ and width $\delta$ of the shear zone (panel d). Notably, $\omega_{0}$ is small even near the rotating base due to slip at the bottom boundary. Moreover, the bulk angular velocity decays rapidly with $h$. The quantitative difference between the profile at $h\,{=}\,10\,\text{mm}$ and those at larger heights is evident in Fig. 2(c). The reduction of angular velocity with increasing $h$ is distinctly nonuniform. Starting from h= 0h\,{=}\,0, the shear zone shifts quickly toward the cylinder axis, accompanied by a rapid decrease in rotation speed. For heights above h= $20\,\text{mm}, h$owever, the rate of change slows markedly. As shown in the inset of Fig. 2(c), the rapid decay of axial angular velocity with $h$ follows a Gaussian trend Cheng et al. (2006); Madani et al. (2021).
Figure 4: (a) Initial configuration of colored tracers placed at mid-height (h=H/2h{=}H{/}2) for two filling levels, covered with a black tracer layer at the top. (b) Reconstructed bulk tracer positions after shear for $H{=}4\,\text{cm}$. (c) Tracer density map in the ($r,h$) plane for $H{=}4\,\text{cm}$ in the steady state, overlaid with the shear zone center and boundaries, indicating a single dominant convection roll. Inset: Side view of the initial arrangement of colored tracers. (d) Quantitative characterization of the vertical positions of the differently colored tracer particles shown in panel (c). Each point denotes the mean color-intensity-weighted position within a vertical strip and the error bars indicate the corresponding standard deviations.
Figure 5: (a) Reconstructed bulk tracer positions after shear for $H{=}2\,\text{cm}$. (b) Individual tracer density maps for $H{=}2\,\text{cm}$ in the steady state, overlaid with the shear zone center and boundaries, indicating two counter-rotating convection rolls. Inset: Side view of the initial arrangement of colored tracers. (c) Tracer density maps for $H{=}2\,\text{cm}$ and bulk heights $h{=}0.5,\,1.0,\,2.0\,\text{cm}$ at an early time point $t\,{=}\,3\,\text{min}$. (d) Mean vertical positions (color-intensity-weighted) of the differently colored tracer particles shown in panel (c).
Comparing $R_{c}$ and $\delta$ in our setup with known results in similar but rough-bottom geometries— where the behavior is well captured by $h{=}H{-}R_{c}\big[1{-}\frac{R_{s}}{R_{c}}\big(1{-}(\frac{H}{R_{s}})^{5/2}\big)\big]^{2/5}$ and $\delta{\sim}h^{\beta}$ ($0.5{<}\beta{<}1$) Fenistein and van Hecke (2003); Unger et al. (2004); Török et al. (2007)— reveals significant deviations in $R_{c}$ and $\delta$ due to slip [Fig. 2(d)]. We confirm this through numerical minimization of energy dissipation using a variational approach Unger et al. (2004); Török et al. (2007); Shaebani et al. (2021); Madani et al. (2021); Unger (2007), showing that reduced bottom friction— characterized by an effective friction $\mu_{\text{rel}}{=}\mu_{\text{bottom}}^{\text{eff}}{/}\mu_{\text{bulk}}^{\text{eff}}$— shifts $R_{c}$ inward [Fig. 2(e)]. As shown in Fig. 2(f), shear zone broadening due to slip enhances with filling height— which increases shear rate asymmetry across the flanks, an essential ingredient for the emergence of convection. These results justify our choice of a smooth-bottom setup to access broader shear zones and thereby amplify the secondary flow signatures.
Surface radial flow.— To probe secondary flow near the surface, we place colored tracer grains (red, blue, green, black) in narrow radial bands at different radii atop the transparent bed [Fig. 3(a)] and track their radial displacement over time. The mean tracer radius $\langle r\rangle$ for each color is monitored at 30-minute intervals for various filling heights $H{=}1{-}4\,\text{cm}$. At $H{=}1\,\text{cm}$, only black tracers near the outer edge ($r{=}85\,\text{mm}$) exhibit slight spreading and other colors remain stationary, indicating negligible surface convection [Fig. 3(b)].
At $H{=}2\,\text{cm}$, tracers at radii $r{=}50,\,75\,\text{mm}$ (near the flanks of the surface shear zone) disperse and converge toward an intermediate radius, while inner tracers at $r{=}15,\,30\,\text{mm}$ remain fixed. This convergence indicates the presence of two counter-rotating convection rolls, featuring either downward flow along the shear flanks and upward return flow between them, or the reverse pattern. During the initial stage (up to $t=3~\mathrm{min}$), black particles at $r=75~\mathrm{mm}$ move outward, while green particles at $r=50~\mathrm{mm}$ move inward. Subsequently, the particles in this region descend through the convection rolls and re-emerge at the surface at radii different from their initial positions, leading to a shift in their mean radii in the opposite direction [Fig. 3(b)]. The initial divergence and subsequent convergence of the black and green particles indicate a downward flow along the shear flanks and an upward return flow between them.
For $H{=}3\,\text{cm}$, the surface flow is strongest: tracers at $r{=}30,\,50\,\text{mm}$ move outward up to radial distance $r{\approx}70\,\text{mm}$, consistent with a dominant single convection roll with downward flow at the outer shear flank at $r{\simeq}70\,\text{mm}$ and upward return flow at the inner shear flank at $r{\simeq}30\,\text{mm}$. Tracers at $r{=}15,\,75\,\text{mm}$ show minimal displacement, indicating they lie outside the active convective region.
At $H{=}4\,\text{cm}$, surface flow is largely suppressed, though there still exists a weak radial motion. The peak radial surface velocity $v_{r}\,{\approx}\,6{\times}10^{-4}\,\text{mm/s}$ is about 2% of the peak azimuthal speed $v_{\theta}\,{\approx}\,3{\times}10^{-2}\,\text{mm/s}$, reflecting the dynamical relevance of convective flows at this height.
Bulk convective flow.— To probe internal convection, we prepare a single colored layer at mid-depth (h=H/2h{=}H{/}2) for two filling heights $H{=}2,\,4\,\text{cm}$. Tracers are arranged in three radial bands: red beads in the inner region ($r{<}6\,\text{cm}$), green in a ring ($6{<}r{<}7\,\text{cm}$), and blue in the outer bulk ($7{<}r{<}10\,\text{cm}$), with a top layer of black tracers covering the free surface [Fig. 4(a)]. After shearing the system for 60 min ($H{=}2\,\text{cm}$) or 180 min ($H{=}4\,\text{cm}$), the system is excavated layer by layer, and the bulk tracer positions are reconstructed [Figs. 4(b) and 5(a)]. For $H{=}4\,\text{cm}$, the tracer density map in the ($r,h$) plane shown in Fig. 4(c,d) reveals a strong single convection roll: red tracers rise while green and blue tracers descend within the shear zone. Black surface tracers are displaced outward and drawn into the bulk at the outer flank. A small surface elevation in the central region suggests dilatancy induced by flow. The entire convective motion is confined within the shear zone, with negligible transport beyond its boundaries.
For $H{=}2\,\text{cm}$, tracer trajectories [Fig. 5(b)] reveal that the flow direction of the two counter-rotating convection rolls is downward at the shear flanks: Tracers from all colors descend along bulk shear flanks, reappear at the top free surface, and move toward the edges of the shear zone, consistent with the surface flow observations. Black surface tracers are drawn toward the surface shear zone edges and disappear into the bulk, further corroborating the counter-rotating rolls. Formation of two counter-rotating convection rolls for $H{=}2\,\text{cm}$ is also evident when we track the colored beads at an earlier time point $t\,{=}\,3\,\text{min}$ for bulk heights $h{=}0.5,\,1.0,\,2.0\,\text{cm}$; see Fig. 5(c,d). Overall, convective transport is localized within the shear zone and highly sensitive to the filling height.
Figure 6: (a) Initial configuration of horizontal red, green, and blue tracer layers at depths $h{=}0.5,\,2,\,3\,\text{cm}$, respectively, in a granular pile of height $H{=}4\,\text{cm}$. (b) Reconstructed positions of the tracers after 160 minutes of shear. (c) 2D tracer density map in the ($r,h$) plane, showing radial and vertical displacements. Shear band center and boundaries are overlaid. Inset: Side view of the initial arrangement of colored tracers. (d) Mean vertical positions (color-intensity-weighted) of the differently colored tracer particles shown in panel (c).
To further improve the bulk resolution at large filling height $H{=}4\,\text{cm}$, we implement a layered tracer configuration. Colored grains are arranged in horizontal planes: red at $h{=}0.5\,\text{cm}$, green at $h{=}2\,\text{cm}$, and blue at $h{=}3\,\text{cm}$ [Fig. 6(a)]. After shearing the system for 160 mins, we excavate the packing layer by layer, capturing the tracer positions at each step [Fig. 6(b)]. This procedure enables us to reconstruct a 2D density map of tracer positions in the ($r,h$) plane [Fig. 6(c,d)]. The tracer displacements reveal clear radial and vertical motions indicative of secondary flow in the bulk. Red particles near the base migrate inward and then upward, tracing a clockwise circulation. Green tracers from the mid-plane predominantly descend and accumulate near the bottom within the shear zone, while a portion at smaller radii is lifted upward— again consistent with toroidal motion. Blue tracers, initially near the top, move downward at larger radii. The overall pattern confirms the presence of a dominant single convection roll confined mainly within the bulk shear zone.
Dilatancy and heap formation.— To quantify surface deformation and dilatancy during shear, we employ a 532.8 nm laser line (Fath Optics) projected at an oblique angle $\alpha\,{=}\,65^{\circ}$ relative to the horizontal, illuminating a diagonal path across the granular free surface [Fig. 7(a)]. Because the transparent glass beads scatter light and broaden the laser sheet, a thin opaque ash layer (thickness < $75\,\mu\text{m}$) is applied to the surface to enhance optical contrast without measurably altering surface roughness. A camera positioned at a fixed angle $\beta\,{=}\,45^{\circ}$ to the horizontal captures images of the illuminated line. For each filling height, six diameters are recorded by rotating the laser-camera assembly around the cylinder axis [see Fig. 7(b)].
From each image, the apparent pixel height $h^{\prime}$ of the illuminated line is extracted using custom MATLAB routines. The true vertical height deviation $\Delta H$ is then obtained through a geometric correction based on the known projection and viewing angles as $\Delta H\,{=}\,h^{\prime}\,\frac{\sin\alpha}{\sin(\alpha{+}\beta)}$. The resulting six profiles are azimuthally averaged to obtain the mean surface-height change, with the standard deviation indicated as shaded envelopes in Fig. 7(c).
The optical resolution of the method depends on the field of view and ranges between 0.06-0.11 mm, sufficient to resolve millimeter-scale heap features given the ${\sim}1$ mm laser-line width. Calibration using reference glass beads of diameter 2 mm yields an accuracy of ± ${\pm}\,0.05$ mm ( ${\sim}\,5\%$ relative error), consistent across filling heights. Tests on smooth reference surfaces with and without the ash layer show $5\%$ variation in measured $\Delta H$, confirming negligible influence from ash-induced micro-roughness.
Figure 7: (a) Schematic of the laser-sheet setup showing the illumination geometry and camera orientation, together with a side-view illustration of the oblique laser projection on the granular surface. (b) Example images of the ash-coated free surface illuminated by the laser line, along with six extracted profiles corresponding to the six measured diameters for the filling height $H\,{=}\,4\,\text{cm}$ (top) and $H\,{=}\,2\,\text{cm}$ (bottom). (c) Azimuthally averaged surface height-change profiles (relative to the initial flat surface) for $H\,{=}\,4\,\text{cm}$ (top) and $H\,{=}\,2\,\text{cm}$ (bottom). Raw digitized data (red points) and smoothed curves (black lines) are shown. (d) Fraction of shear rates at the outer and inner flanks of the surface shear zone vs $H$. The asymmetry grows with filling height, promoting a transition from symmetric counter-rotating convection rolls to a dominant single roll as shear becomes localized more strongly at the outer edge.
For clarity of presentation, the raw profiles (red points) are shown together with smoothed curves obtained via robust LOESS filtering (span 0.04 for $H\,{=}\,2\,\text{cm}$ and 0.08 for $H\,{=}\,4\,\text{cm}$), which suppress digitization noise and grain-scale curvature while preserving physically meaningful variations. As shown in Fig. 7(c) for $H\,{=}\,2\,\text{cm}$, a pronounced elevation develops within the shear zone ($r\,{\approx}\,50{-}75$ mm), reaching a maximum $\Delta H\,{\approx}\,2.2\,\text{mm}$ (${\sim}11\%$ of $H$). For $H\,{=}\,4\,\text{cm}$, the surface exhibits a broader central rise of $\Delta H\,{\approx}\,2.5\,\text{mm}$ (${\sim}7\%$ of $H$) within the shear zone. Although these dilatancy-induced height changes are modest compared with the large heaps commonly reported for elongated particles Fischer et al. (2016); Wortel et al. (2015), their magnitude and radial position are fully consistent with the emergence of a single roll or a pair of counter-rotating convection rolls at different filling heights.
Onset of convection roll transition.— To rationalize the transition from two counter-rotating convection rolls at low filling heights to a single dominant roll at larger fillings, we examine how the radial shear-rate distribution evolves with the broadening of the shear zone. Using the error-function form of the surface angular-velocity profile $\omega(r)$, the associated tangential velocity is $v_{\theta}(r){=}\,r\,\omega(r)$, from which the radial shear rate is obtained as $\dot{\gamma}(r){=}\frac{1}{2}(\partial_{r}v_{\theta}{-}v_{\theta}{/}r)=\frac{1}{2}r\partial_{r}\omega$. The shear-rate profile exhibits two peaks located around the flanks of the shear zone (defined as $R_{c}{\pm}\delta$). We estimate the shear rates at these two locations and plot their ratio $\dot{\gamma}_{\text{out}}{/}\dot{\gamma}_{\text{in}}$ as a function of filling height $H{\propto}\,\delta^{1{/}\beta}$ (with $\beta{=}0.5$) in Fig. 7(d). This ratio increases monotonically with $H$, indicating a growing shear-rate imbalance between the outer and inner flanks.
To connect this shear-rate asymmetry to the emergence of secondary convection, we assume a monotonic relation between the local shear rate and local dilation Khosropour et al. (1997). Thus, large dilatancies occur around the two peaks of shear rate located at the flanks of the shear zone. At small filling heights, the two shear-rate peaks are nearly symmetric, and the corresponding dilatancy regions, where the material locally expands and becomes more mobile, are of comparable strength. Because gravity biases particle motion downward in locally dilated regions, both flanks support similarly strong downward flow components, resulting in the coexistence of two counter-rotating convection rolls. As the filling height increases, the shear zone widens and the shear-rate profile becomes increasingly asymmetric: the outer flank experiences significantly stronger shear, producing a stronger downward driving component for convection. Enhanced dilatancy reduces local packing fraction and facilitates particle rearrangements, allowing gravity-driven downward motion to dominate on the outer side. The inner flank, experiencing weaker shear and weaker dilatancy, can no longer sustain an opposing convective roll. This imbalance naturally yields the experimentally observed transition to a single, outer-driven roll at larger filling heights. Our measured surface-height profiles shown in Fig. 7(c) are consistent with this model; the profiles reveal dilatancy-induced elevations whose magnitude and radial location correlate with the proposed single roll or two counter-rotating convection rolls for H= 4H\,{=}\,4 or $2\,\text{cm}$.
This qualitative framework, linking the filling-height-dependent imbalance between the two shear-rate peaks with gravity-biased motion in locally dilated regions, provides a consistent minimal mechanism for both (i) the formation of secondary convection rolls in spherical granular packings and (ii) their transition from two rolls to a single roll as $H$ increases. It also suggests that similar transitions might occur in tall granular piles. At large depths, however, gravity-induced confining stresses suppress particle rearrangements Shaebani, Unger, and Kertész (2007, 2008) and thereby inhibit the formation of convection rolls. Under reduced-gravity conditions Murdoch et al. (2013); Born, Schmitz, and Sperl (2017), this constraint is partially relaxed. Yet a lower gravitational acceleration also weakens the gravity-driven downward motion that is essential for sustaining convection. One may further argue that reduced gravity diminishes frictional interparticle forces, leading to reduced dilatancy and thus an additional weakening of secondary convection. Consistent with these expectations, experiments have indeed reported suppressed convective flows under microgravity Murdoch et al. (2013).
We note that granular flows in the quasistatic regime, as studied here, are fully reproducible. For a given filling height and shear protocol, both the primary flow field and the secondary convection pattern remain robust and do not vary between experimental runs. The error bars presented in our analysis therefore reflect intrinsic tracer dispersion, tracer diffusion, and finite particle-size effects, as the dominant sources of experimental variability in our system.
We have presented experimental evidence of geometry-driven convective flow in sheared granular packings of spherical particles. Using tracer layers and tomographic reconstruction, we reveal a transition from symmetric counter-rotating convection rolls at low filling heights to a dominant single-roll pattern at higher fillings. This transition correlates with increasing asymmetry in the radial shear profile, consistent with a shear-driven convection mechanism. Our work provides direct experimental evidence that secondary convection rolls can arise in the bulk of spherical granular materials, even in the absence of particle-shape anisotropy or boundary-driven effects, extending beyond previous simulation-based studies. Such geometry-induced convection can play a crucial role in mixing, segregation, and particle migration in granular media, with direct relevance to natural processes and industrial applications.
DATA AVAILABILITY
The data that support the findings of this study are available from the corresponding author upon reasonable request.